Please help...physics
Estimate the cost of cooking a turkey for 4hours in an oven that operates continuously at 20A and 240V.
If the energy is purchased at an estimated price of 8.00$ per kilowatt hour, the cost is?

Answer by rothauserc(4718)

Energy(kWh) = Power(kW) * hours
:
Power(kW) = Amps * voltage/1000
:
we are given 4 hours at 20 amps and 220 volts with 8.00$/kWh
:
Power(kW) = 20 * 240 / 1000 = 4.8 kW
:
Energy(kWh) = 4.8 * 4 = 19.2 kWh
:
Cost = 8.00$ * 19.2 = 153.6$
: